OpenStudy (cookie_2046):

Which is 216 written as a power of a number? A. 63 B. 36 C. 1082 D. 723 i think a :P

OpenStudy (usukidoll):

\[4a^3-2b^2+5\] we could just plug in those values given and solve

OpenStudy (usukidoll):

seriously?

OpenStudy (cookie_2046):

Something wrong?

OpenStudy (usukidoll):

\[4(3)^3-2(-2)^2+5\] \[4(27)-2(4)+5 = 108-8+5=105\]

OpenStudy (usukidoll):

that 3 and 2 next to that variable should have be a to the third power and b to the second power. unless the problem was really written like that.

OpenStudy (usukidoll):

scan the original problem right here right now . so we can straighten this out

OpenStudy (cookie_2046):

@kidrah69 do u know the answer?

OpenStudy (usukidoll):

I swear the equation is \[4a^3-2b^2+5 \]

OpenStudy (usukidoll):

and you just plug in a = 3 and b = -2 ... do the simple arithmetic.. that's it.

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @UsukiDoll I swear the equation is \[4a^3-2b^2+5 \] \(\color{blue}{\text{End of Quote}}\) yep right

OpenStudy (usukidoll):

THANK YOU!

Nnesha (nnesha):

LO_OL

OpenStudy (usukidoll):

how the op got over 70,000 as the answer is beyond me ... aww it got deleted. but the guess was 79501 or something like that... I was like WHATTTTTTTTTTTTTT

OpenStudy (usukidoll):

OMG HE EDITED THE QUESTION TOO WHAT THE

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @UsukiDoll OMG HE EDITED THE QUESTION TOO WHAT THE \(\color{blue}{\text{End of Quote}}\) ohh now i got it

OpenStudy (cookie_2046):

umm i'm gonna go .-.

Nnesha (nnesha):

too bad cO_okie stop asking same question again and again

OpenStudy (cookie_2046):

what r u talking about?

OpenStudy (usukidoll):

you keep on asking the same thing over and over again and tried to delete your evidence and make me and Nnesha look bad. we saw right through it...